OpenStudy (anonymous):

Decide which of the following are solutions to this system of equations: y = 2x + 3 and y = x2 - 5 a. (4,11) b. (1,-4) c. (-3,-1) d. (-2,-1)

OpenStudy (anonymous):

How would you solve the equation ?

OpenStudy (anonymous):

the answer choice that I have is both a and c

OpenStudy (anonymous):

@Koikkara

OpenStudy (anonymous):

@Owlcoffee

OpenStudy (owlcoffee):

\[y=2x+3\] \[y=x^2-5\] Since both equations are equal "y", must mean that both left sides of each equation must have the same values, so all we do is replace the given equations: \[x^2-5=2x+3\] And order it up to the general form: \[x^2-2x-8=0\] And now, all you have to do is solve that quadratic equation using the general formula.

OpenStudy (anonymous):

How many solutions are there to the system: \begin{array}{l}y = 1 + 2x + {x^2}\\y = 1 - x\end{array} Write your answer as an integer. could you help me with this one I have 2 for the answer

OpenStudy (owlcoffee):

The amount of solutions are always determined by the variable of highest exponent: we wil ldo the same as we did previously: \[1-x=1+2x+x^2\] and order it: \[x^2+3x=0\] Since there is a x^2 it must mean that the system has two solutions.

OpenStudy (anonymous):

Find the solution to the system using a table. Write your answers as ordered pairs using parentheses. If there are two solutions, separate the two solutions with a comma. Do not put a space between them. y = 3 y = 3x this one I have (0,0) for the answer

OpenStudy (owlcoffee):

Well, on that excercise you have to draw two columns and then assign up to three values to "x", let them be 0,1 and 2. The solution will be the value of "x" that has the same value on both functions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

would it be ),1 and 0,3

OpenStudy (anonymous):

0,1